Problem 2.32 - Enhanced- with Feedback Review| Constants | Periodic Table PartA What is the final speed? Express your answer with the appropriate units. Light-rail passenger trains that provide transportation within and between cities speed up and slow down with a nearly constant (and quite modest) acceleration. A train travels through a congested part of town at 60 m/s. Once free of this area, it speeds up to 11 m/s in 8.0 s. At the edge of town, the driver again accelerates, with the same acceleration, for another 16 s to reach a higher cruising speed 図? vValue Units de Previous
Answer #1

acceleration of the train,

a = ( v - u) / t

a = (11-6)/ 8

a = 0.625 m/s^2

now again

v = u + at

v = 11 + 0.625* 16

v = 21 m/s

==========

do comment in case any doubt, will reply for sure . Goodluck